A gas at pressure P1 N/m2 and temperature 30°C is heated to 61° C at constant volume. Find its new pressure
1
Expert's answer
2019-04-05T10:27:27-0400

At constant volume:


"\frac{p}{T}=const"

"\frac{p_1}{T_1}=\frac{p_2}{T_2}"

"p_2=p_1 \frac{T_2}{T_1}=p_1 \frac{61+273}{30+273}=1.1p_1"